Apex Company invested $14,400 in new equipment. The more efficient new equipment was expected to reduce operating cash outflows over the next five years by the following:

yr1 = $8000 yr2= $6000 yr3= $5000 yr4= $6000 yr5= $5000

Using the averaging method, the payback period for the investment in the oven would be:
A) 3.0 years B) 2.4 years C) 2.0 years D) 1.7 years

Please show work and no excel sheets. Thanks
Answer #1

Answer: Option (B) 2.4 years

Annual cash flow = 8000 + 6000 + 5000 + 6000 + 5000 = 30000

= 30000 / 5

Annual cash flow = 6000